You need to use your cell phone, which broadcasts an 750 MHz signal, but you're behind two massive, radio-wave-absorbing buildings that have only a 19 m space between them. What is the angular width of the electromagnetic wave after it emerges from between the buildings?
Answer #1

dsin\theta = n\lambda

\lambda = c/f

d = 19 m , n =1 , c =3*10^8 m/s, f =750 MHz

19sin\theta = 1*3*10^8/(750*10^6)

\theta = 1.21 degrees

angular width = 1.21*2 = 2.42 degrees